How many ml each of a 15% acetic acid solution and a 30% acetic acid solution must be mixed to obtain 600ml of a 22% acetic acid solution?
Answer by stanbon(75887)

How many ml each of a 15% acetic acid solution and a 30% acetic acid solution must be mixed to obtain 600ml of a 22% acetic acid solution?
----------
Equation:
acid + acid = acid
0.15x + 0.30(600-x) = 0.22*600
-----
15x + 30*600 - 30x = 22*600
------
-15x = -8*600
x = 320 mL (amt. of 15% solution needed)
------
600-x = 280 mL (amt. of 30% solution needed)
------------
Cheers,
Stan H.
==================